# Thermal Stoner-Wohlfarth Model for Magnetodynamics of Single Domain Nanoparticles: Implementation and Validation

### v1

@@ II.1 Static Properties
At a finite temperature, an ensemble of magnetic colloids in a static homogeneous magnetic field is in a thermodynamic equilibrium at times larger than the systems characteristic relaxation time. For any particular configuration of anisotropy axes in a nanoparticle ensemble, the equilibrium state is uniquely defined by ($\sigma$, $\xi$). In Fig. 4, we scrutinise the tSW model in its ability to correctly reproduce equilibrium magnetisation, $m$, of an ensemble of magnetic nanoparticles. The equilibrium magnetization of the system is the total sum of the magnetic moments $\mu$, projected onto the axis of the applied magnetic field. As previously mentioned, we distinguish cases where the particles exhibit either only Néel (solid superparamagnetic) or both Brownian and Néel (non-interacting ferrofluid) relaxation mechanisms.
Within the tSW model, the discrete states in which the magnetic moment can be for a solid superparamagnet, correspond to the energy minima in the total magnetic energy (Eq. 1). The equilibrium magnetisation can be obtained if one performs a summation over a set of allowed discrete orientations:
@@
The problem of describing the equilibrium magnetisation of an ensemble of immobilized, non-interacting, randomly oriented uniaxial particles, has been solved based on a generalization of the Langevin function for the case of solid dispersions with random orientation texture [60, 61, 62]. The full expression together with an outline of the derivation are provided in the Appendix A.
In the case of a non-interacting ferrofluid, $\vec{n}$ and $\vec{e}$ are no longer independent variables - in the tSW model, there exists a unique set of allowed magnetic moment orientations, for any given orientation of the anisotropy axis. The equilibrium magnetisation can be obtained in a similar form to Eq. 15, where:
@@
For $\xi=0$, the distribution is random and $f_{n}^{tSW}=1$. In general, however, the magnetization of a non-interacting ferrofluid is given by the Langevin function [34], and does not depend on the anisotropy parameter $\sigma$.
@@
Looking at Fig. 4, one can see that the absolute magnetisation error $\Delta m=|m^{tSW}-m^{sim}|$ demonstrates that the tSW model can reproduce the equilibrium magnetisation within a few percent of the ground truth, in all cases except for particle sizes and/or temperatures where $\sigma\leq 5$, where one can expect a purely super-paramagnetic response. Below this $\sigma$ range, we encounter the limitations of the model. Let us address these limitations. In a solid superparamagnet, the only relaxation mechanism that the particles can leverage to reach equilibrium is Néel.
@@
In the tSW model, there is only a single state in which the dipole moment can be above $h_{cr}$. In other words, the tSW model is equivalent to the classical SW model above $h_{cr}$. In general, however, the magnetic moment orientation can fluctuate relative to the states allowed in the classical SW model [60, 61, 62]. The frequency of these intrawell fluctuations is inversely proportional to $\sigma$. With this in mind, one can understand why the tSW model overestimates the equilibrium magnetisation of low $\sigma$ particles in the large $\xi$ region. Moreover, the intrawell modes matter for high $\xi$ regardless of $\sigma$, as Néel relaxation ($1/p_{1}$) is no longer sufficient for an accurate description of magnetic relaxation [63, 36]. Hence, we observe regions with a slight increase in $\Delta m$ with growing $\xi$. Finally, the importance of intrawell states relative to interwell states increases with decreasing $\xi$, since the applied field is not strong enough to constrain the dipole moment fluctuations perpendicular to the anisotropy axis.
@@
Moving on to the non-interacting ferrofluid system, where both Néel and Brownian relaxation are at play, the same conclusions largely hold, as for a solid superparamagnet. However, one can note a marginally worse agreement between the tSW model and the continuum-level description. Consider the high $\xi$ region, where one can assume Néel relaxation plays a minimal role. For high $\sigma$, we have seen that for a fixed random distribution of the anisotropy axis, the SW reproduces the equilibrium state very well. In the same circumstances, given a non-random distribution of anisotropy axis, the relative error is almost quadrupled. The relative error being different between the left and right subplot in Fig. 4, is related to the fact that, for a solid superparamagnet, the anisotropy axis orientations are randomly distributed and fixed, whereas for a ferrofluid this is not the case. The classical SW model estimates the equilibrium magnetisation incorrectly, but not equally incorrectly for all anisotropy axis distributions. In fact, it can be inferred by comparing the left and right side subplots in Fig. 4, that the more the anisotropy axes are aligned with the magnetic field, the worse is the equilibrium state estimation in the classical SW model.
@@
Finally, it is worth highlighting a particularity of the tSW model, related to the initial susceptibility for low $\sigma$. Initial magnetic susceptibility is an important characteristic of the material linear response to a weak applied field (or “probing field”):
$\chi=\frac{\partial m}{\partial H}\Big{|}_{H\rightarrow 0}.$
@@
$\chi^{tSW}(0)=\chi_{L}(1+1/\sigma).$ (18)
@@
In the limit of no particle anisotropy, the static susceptibility should correspond to the Langevin susceptibility. However, this is not the case in the tSW model, where the values diverge. Although the tSW model reproduces the characteristic internal relaxation time related with Néel relaxation well, it is inadequate for simulations of small enough nanoparticles, or at high enough temperatures where $\sigma<5$.
@@
Having addressed the limitation of the tSW at length, it is important to note that the absolute error for $\sigma\geq 5$ remains less than $10\%$, which qualifies the tSW approach as surprisingly accurate within a wide ($\sigma,\xi$) range of applicability. It is apparent that the dominant internal relaxation mechanism is the Néel one, even for relatively small $\sigma$. Furthermore, our implementation of the tSW is in agreement with the theoretical prediction for the model across the board. The key insight here is that the validity of the tSW model is most constrained by the limitations of the classical SW model, which are well understood and documented. The discrepancies stemming from the lack of intrawell modes are less important, and we do not see this as a substantial additional limitation to the model. Moreover, as long as one is considering non-interacting systems, one can compensate for the overestimation stemming from the classical SW model by tuning the saturation magnetisation of the colloids to fit the correct equilibrium magnetisation curves.
@@
In summary, the tSW model is a reasonably accurate approach to introduce internal magnetisation dynamics for $\sigma\geq 5$ nanoparticles via Néel relaxation, that can reproduce the correct equilibrium in this $\sigma$ range. Having established that, the discussion can proceed to the study of dynamics, which is the type of study where the tSW model can benefit researchers interested in long timescales and bulk sized systems.
@@ II.2 Dynamic Properties
One of the key collective properties of any magnetic soft matter system is its characteristic magnetic relaxation time. This is especially relevant for systems designed for heating applications. The magnetic relaxation time can be assessed by calculating the dynamic susceptibility spectra and locating the maximum/maxima of the susceptibilities imaginary part.
Figure 5: Dynamic susceptibilities for a solid superparamagnet (left column) and non-interacting ferrofluid (right column) as a function of $\omega\tau_{D}$, where $\tau_{D}/\tau_{B}=0.02$. The real and imaginary parts of the susceptibility are shown separately (top and bottom rows, respectively) for clarity. In each subplot, profiles for a continuum description, theoretical and simulation tSW model results are superimposed (see legend), for different values of $\sigma$ (see colourbar).
Let us consider an ensemble of immobilized, non-interacting, randomly oriented particles with uniaxial magnetic anisotropy. In a harmonically oscillating field $H=H_{0}\cos\omega t$, where $t$ is the time and $\omega$ is the oscillation frequency, the magnetisation can generally be represented as a series:
$m(t)=\sum_{k=1}^{\infty}\left(m^{\prime}_{k}(\omega)\cos k\omega t+m^{\prime\prime}_{k}(\omega)\sin k\omega t\right),$
@@
$\chi^{\prime}(\omega)=m^{\prime}_{1}(\omega)/H_{0},\>\>\chi^{\prime\prime}(\omega)=m^{\prime\prime}_{1}(\omega)/H_{0}.$
@@
$\chi=\left(\frac{1}{3}\chi_{\parallel}+\frac{2}{3}\chi_{\perp}\right)$
This relation must be true for both the static and dynamic case. For immobilized particles, within the continuous approach and without gyromagnetic effects [52]:
$\chi_{\parallel}(\omega)\simeq\frac{\chi_{\parallel}(0)}{1-i\omega\tau_{\parallel}},\>\>\>\chi_{\parallel}(0)=\chi_{L}(1+2S),$ $\>\>\>\tau_{\parallel}=\tau_{D}\frac{e^{\sigma}-1}{2\sigma}\left[\frac{1}{1+1/\sigma}\sqrt{\frac{\sigma}{\pi}}+2^{-\sigma-1}\right]^{-1};$ (20)
$\chi_{\perp}(\omega)\simeq\frac{\chi_{\perp}(0)}{1-i\omega\tau_{\perp}},\>\>\>\chi_{\perp}(0)=$ $\chi_{L}(1-S),\>\>\>\tau_{\perp}=2\tau_{D}\frac{1-S}{2+S},$
where the order parameter $S$ is given by:
@@
$D(x)=e^{-x^{2}}\int_{0}^{x}e^{t^{2}}\,\text{d}t,$
and $\chi_{L}$ is the static susceptibility of an isotropic superparamagnet (Langevin susceptibility):
$\chi_{L}=\frac{\mu_{0}\mu^{2}N}{3Vk_{B}T}.$ (22)
In the case where particles are not immobilised, it is sufficient to replace relaxation times with effective values [64]:
$\tau^{*}_{\parallel}=\left(\tau^{-1}_{\parallel}+\tau_{B}^{-1}\right)^{-1},\\ \tau^{*}_{\perp}=\left(\tau^{-1}_{\perp}+\tau_{B}^{-1}\right)^{-1}.$
In the tSW model, the relaxation time $\tau_{\parallel}$ is an input parameter, which means that the dynamics of tSW parallel to the easy axis must be the same as in the continuous case. The orthogonal relaxation time $\tau_{\perp}$ is instantaneous. Therefore, the dynamic susceptibility of an ensemble of frozen, uniaxial magnetic nanoparticles with a random distribution of anisotropy axis is given by:
@@
Once again, for a liquid, it is sufficient to introduce effective relaxation times: $\tau_{\parallel}\rightarrow\tau^{*}_{\parallel}$. Note, that the orthogonal magnetic response (which is instantaneous and does not have any fluctuations associated with it) will not make any contribution to the zero-field magnetization autocorrelation function. Therefore, the calculated susceptibilities from the autocorrelation function when using the tSW method will be incorrect. The simulated susceptibilities have been calculated by applying a weak oscillating magnetic field (maximum amplitude corresponding to $\xi=0.1$) for a span of frequencies, in order to be able to determine the complex susceptibilities from the time evolution of the total magnetisation vector (see Eq.19). Similarly, the static susceptibilities should be estimated from the initial slope in the magnetisation profiles.
@@
Looking at Fig. 5, we can see that the tSW model reproduces the correct dynamics in both the solid superparamagnet and non-interacting ferrofluid systems, as quantified by the initial dynamic susceptibilities. Of course, in the initial susceptibility regime, the tSW model incorporated dipole moment fluctuations only along the anisotrpy axis associated with $\tau_{\parallel}$, which in this case is equivalent to $\tau_{N}$. Since this is an input parameter for the model, the tSW can confidently reproduce the correct internal dynamics for the lowest frequency relaxation process ($\tau_{\parallel}$). It is apparent that Néel relaxation is very relevant even for $\sigma=10$ particles, where the dynamic susceptibility is clearly distinct from the Debye limit shown with dots. Recall that the maximum of the Debye limit is at $\omega\tau_{d}=0.01$, which is set by the model as $\tau_{D}/\tau_{N}=0.01$. The high frequency peaks, resolved in the continuum description, corresponding to $\tau_{\perp}$, cannot be captured in the tSW, simply due to the fact that there are no fluctuations associated perpendicular to the anisotropy axis in the $H\rightarrow 0$ regime. Outside of the initial susceptibility regime, this is not an issue as the dipole moment instantaneously follows the local energy minima, as obtained in the classical SW model, where there is always a relative angle the dipole moment takes with respect to the anisotropy axis. However, within the range of applicability we suggest ($\sigma>5$), these higher frequency processes are distinct enough that they could be averaged out while still resolving the correct dynamics. Having said that, it should also be clear that the in-field susceptibility contribution perpendicular to the anisotropy axis is frequency independent.
@@
As suggested at the beginning of this section, the position(s) of the maxima of the imaginary part of the initial susceptibility indicate the optimal field frequencies for heating applications. Fig. 5 clearly shows that if one needs to optimise the frequency of an applied magnetic field for hyperthermia depending on the particle size and material, using a point dipole approximation can be very misleading, as even for rather high anisotropy energies, the shape and the imaginary part of the susceptibility spectra drastically differ from that of a purely Brownian, or purely Neél relaxing systems.
@@
In summary, we can see that the tSW approach can confidently be used to incorporate Néel relaxation and simulate the dynamics associated with that relaxation process for magnetic nanoparticles with $\sigma\geq 5$. This is also the first major use case where we see that this model could enable new simulation studies. The price to pay in terms of accuracy is slight, compared to the simulation scale compromises that would be unavoidable with an approach that reproduces also the higher frequency relaxation processes. The second major use case of the tSW model would be to simulate systems where $\tau_{B}$ is high. In this case, a Brownian dynamics approach to thermalise each particle might be impractical. A good example here would be multicore or composite magnetic colloids, where the nanoparticles are embedded in a solid matrix[65, 66, 67].

### v2

@@ II.1 Static Properties
At a finite temperature, an ensemble of magnetic colloids in a static homogeneous magnetic field reaches thermodynamic equilibrium at times larger than the system’s characteristic relaxation time. For any particular configuration of anisotropy axes in a nanoparticle ensemble, the equilibrium state is uniquely defined by ($\sigma$, $\xi$). In Fig. 4, we scrutinize the tSW model in its ability to correctly reproduce equilibrium magnetization, $m$, of an ensemble of magnetic nanoparticles. The equilibrium magnetization of the system is the total sum of the magnetic moments $\mu$, projected onto the axis of the applied magnetic field, and normalized to unity throughout the manuscript. As previously mentioned, we distinguish cases where the particles exhibit either only Néel (solid superparamagnet) or both Brownian and Néel (ferrofluid) relaxation mechanisms.
Within the tSW model, the discrete states in which the magnetic moment can be for a solid superparamagnet, correspond to the energy minima in the total magnetic energy (Eq. 1). The equilibrium magnetization can be obtained if one performs a summation over a set of allowed discrete orientations:
@@
The problem of describing the equilibrium magnetization of an ensemble of immobilized, non-interacting, randomly oriented uniaxial particles, has been solved based on a generalization of the Langevin function for the case of solid dispersions with random orientation texture [70, 71, 72]. The full expression together with an outline of the derivation are provided in the Appendix A.
In the case of a non-interacting ferrofluid, $\vec{n}$ and $\vec{e}$ are no longer independent variables - in the tSW model, there exists a unique set of allowed magnetic moment orientations, for any given orientation of the anisotropy axis. The equilibrium magnetization can be obtained in a similar form to Eq. 15, where:
@@
For $\xi=0$, the distribution is random and $f_{n}^{tSW}=1$. In general, however, the magnetization of a non-interacting ferrofluid is given by the Langevin function [34]:
$m^{L}=mL(\xi),\ \ L(\xi)=\coth(\xi)-\frac{1}{\xi}.$ (18)
In this case, the magnetization does not depend on the anisotropy parameter $\sigma$.
@@
In Fig. 4, we are showing the absolute magnetization error, $\Delta m=100\cdot|m^{tSW}-m^{cont}|/m$, as a function of $\xi$. Note that $m^{tSW}$ represents either the simulated data (symbols) or the analytical results based on the tSW model (lines). Specifically, $m^{tSW}$ corresponds to Eq. 15 for a solid superparamagnet and Eq. 17 for a ferrofluid. Similarly, $m^{cont}$, representing the continuum-level description, is given by Eq. 45 for a solid superparamagnet and Eq. 18 for a ferrofluid. Fig. 4 demonstrates that the tSW model can reproduce the equilibrium magnetization within a few percent of the ground truth, in all cases except for particle sizes and/or temperatures where $\sigma\leq 5$, where one can expect a purely super-paramagnetic response. Below this $\sigma$ range, we encounter the limitations of the model. Let us address these limitations. In a solid superparamagnet, the only relaxation mechanism that the particles can leverage to reach equilibrium is Néel.
@@
In the tSW model, there is only a single state in which the dipole moment can be above $h_{cr}$. In other words, the tSW model is equivalent to the classical SW model above $h_{cr}$. In general, however, the magnetic moment orientation can fluctuate relative to the states allowed in the classical SW model [70, 71, 72]. The frequency of these intrawell fluctuations is inversely proportional to $\sigma$. With this in mind, one can understand why the tSW model overestimates the equilibrium magnetization of low $\sigma$ particles in the large $\xi$ region. Moreover, the intrawell modes matter for high $\xi$ regardless of $\sigma$, as Néel relaxation ($1/p_{1}$) is no longer sufficient for an accurate description of magnetic relaxation [73, 36]. Hence, we observe regions with a slight increase in $\Delta m$ with growing $\xi$. Finally, the importance of intrawell states relative to interwell states increases with decreasing $\xi$, since the applied field is not strong enough to constrain the dipole moment fluctuations perpendicular to the anisotropy axis.
@@
Moving on to the non-interacting ferrofluid system, where both Néel and Brownian relaxation are at play, the same conclusions largely hold as for a solid superparamagnet. However, one can note a marginally worse agreement between the tSW model and the continuum-level description. Consider the high $\xi$ region, where one can assume Néel relaxation plays a minimal role. For high $\sigma$, we have seen that for a fixed random distribution of the anisotropy axis, the SW reproduces the equilibrium state very well. In the same circumstances, given a non-random distribution of anisotropy axes, the relative error is almost quadrupled. The difference in relative error between the left and right subplots in Fig. 4 is related to the fact that, for a solid superparamagnet, the anisotropy axis orientations are randomly distributed and fixed, whereas for a ferrofluid, this is not the case. The classical SW model estimates the equilibrium magnetization incorrectly, but not equally incorrectly for all anisotropy axis distributions. In fact, it can be inferred by comparing the left and right side subplots in Fig. 4, that the more the anisotropy axes are aligned with the magnetic field, the worse is the equilibrium state estimation in the classical SW model.
@@
Finally, it is worth highlighting a particularity of the tSW model, related to the initial susceptibility for low $\sigma$. Initial magnetic susceptibility is an important characteristic of the material’s linear response to a weak applied field (or “probing field”):
$\chi=\frac{\partial m}{\partial H}\Big{|}_{H\rightarrow 0}.$ (19)
@@
$\chi^{tSW}(0)=\chi_{L}(1+1/\sigma),$ (20)
where $\chi_{L}$ is the static susceptibility of an isotropic superparamagnet (Langevin susceptibility):
$\chi_{L}=\frac{\mu_{0}\mu^{2}N}{3Vk_{B}T}.$ (21)
@@
In the limit of no particle anisotropy, the static susceptibility should correspond to the Langevin susceptibility [34]. However, this is not the case in the tSW model, where the values diverge. Although the tSW model reproduces the characteristic internal relaxation time related with Néel relaxation well, it is inadequate for simulations of small enough nanoparticles, or at high enough temperatures where $\sigma<5$.
@@
Having addressed the limitation of the tSW at length, it is important to note that the absolute error for $\sigma\geq 5$ remains less than $10\%$, which qualifies the tSW approach as surprisingly accurate within a wide ($\sigma,\xi$) range of applicability. It is apparent that the dominant internal relaxation mechanism is the Néel one, even for relatively small $\sigma$. Furthermore, our implementation of the tSW is in agreement with the analytical prediction for the model across the board. The key insight here is that the validity of the tSW model is most constrained by the limitations of the classical SW model, which are well understood and documented. The discrepancies stemming from the lack of intrawell modes are less important, and we do not see this as a substantial additional limitation to the model. Moreover, as long as one is considering non-interacting systems, one can compensate for the overestimation stemming from the classical SW model by tuning the saturation magnetization of the colloids to fit the correct equilibrium magnetization curves.
Figure 5: The magnetization $m$ as a function of $\xi$, for a solid superparamagnet (orange for $\chi_{L}=2.4$, red for for $\chi_{L}=4.8$) and a ferrofluid (violet). Interacting systems are represented by square symbols (simulation data) and dashed lines (theory), while non-interacting systems use circles (simulation data) and dotted lines (theory). Dotted lines correspond to Eqs. 17 and 15 (ferrofluid and solid superparamagnet, respectively). Dashed lines represent MMTF extensions of these equations, with an additional scaling factor $\alpha(\xi)=a+b\xi$ for the solid superparamagnet (details in text) where $(a,b)$ are fitted parameters (for $\chi_{L}=2.4$, $a=0.8151$ and $b=0.0213$; $\chi_{L}=4.8$, $a=0.4989$ and $b=0.0778$). Interacting systems were simulated at $\lambda=3$. The effects of demagnetization fields in have been removed from the simulation data.
Figure 6: Visualization of the dipole moment orientations in: (a) non-interacting; (b) interacting solid superparamagnet ($\chi_{L}=4.8$), at $\xi=5$. The magnetic field is applied along the $z$-axis. Color-coding corresponds to the $z$-component of the dipole moment, as indicated by the color bar. The arrow size of the dipole moments is exaggerated for visualization purposes and has no physical meaning.
The tSW model can also be used to study interacting systems. From the point of view of the model, there is no difference between interacting and non-interacting systems. The model requires a magnetic field as an input parameter. If one is studying interacting systems, it is necessary to pass an effective field (a vectorial sum of the total dipole field the nanoparticle feels and the applied magnetic field) to the solver, instead of the applied magnetic field. All other aspects of dipole-dipole interactions are exogenous to the model and must be handled separately. In Fig.5, we consider how dipole-dipole interactions affect the equilibrium magnetization for $\sigma=5$ solid superparamagnet and ferrofluid systems, respectively. For the parameters we used, a $\sigma=5$ nanoparticle corresponds to the dipolar coupling parameter $\lambda=3$. The $\lambda$ parameter quantifies the strength of the dipole-dipole interaction with respect to thermal fluctuations, when two dipoles are collinear and at a touching distance, and is given by:
$\lambda=\frac{\mu_{0}\mu^{2}}{4\pi d^{3}k_{b}T}.$ (22)
While the magnetization of non-interacting ferrofluids is known to follow the Eq. 18, the magnetization of interacting ferrofluids is well described with the modified mean-field theory (MMFT) instead, originally presented in Ivanov and Kuznetsova [74], as long as dipolar forces do not lead to significant cluster formation [75]. The first-order MMFT expansion of the Langevin parameter $\xi_{e}=\xi+\chi_{L}m^{L}(\xi)$, leads to the following magnetization: $m=m^{L}(\xi_{e})$. We applied an MMFT correction to Eq.17, to obtain an analytical expression of an interacting ferrofluid within the scope of the tSW model (purple dashed line). This simple treatment is in good agreement with our simulated data for an interacting ferrofluid (purple square symbols). We reproduce the expected behavior, which is that dipole-dipole interactions tend to increase the magnetic response in a ferrofluid. However, it must be said that it is a non-trivial claim that the MMFT approach should work for a two-state model such as the tSW model, and a more rigorous derivation would be necessary. Comparing the simulated data for non-interacting and interacting solid superparamagnet systems at $\chi_{L}=2.4$ (orange circles and square symbols, respectively), we see that interactions lead to only a slight increase in the magnetic response, that differs qualitatively from what one might expect from Elfimova et al. [34]. For $\chi_{L}=4.8$ (red square symbols), we can note pronounced decrease in magnetization in the low to moderately high $\xi$ range. One can get an intuitive sense of what could be the mechanism behind these results by looking at Fig. 6. Comparing 6(a) and 6(b), it should be visually accessible that the dipoles in a non-interacting system are more aligned with the applied field than for an interacting system. In an interacting system, the dipoles are coupling to a local field, which is a complex function of the dipolar fields and the applied field. In the case of an interacting solid superparamagnet, our conjecture is that dipole-dipole interaction correlated vortexes are formed, which impedes the magnetic response of the system. The mechanism is reminiscent of the one of ring formation in strongly interacting magnetic fluids [76] and vortex formation in magnetic multicore particles [77, 78]. In analogy to these works, we consider these vortices as ”magnetic holes” that can be effectively subtracted from the system’s magnetic response. The functional dependence of the fraction of magnetic holes on the applied magnetic field strength is unknown and a matter for a separate, in-depth study. For the purposes of this work, we assumed that the fraction of magnetic holes decreases linearly with an applied magnetic field strength. With this assumption, we can incorporate dipole interactions using an MMFT based extension to Eq. 15 to obtain an analytical form for the magnetization for an interacting solid superparamagnet, shown in Fig.5, in the following form: $m^{tSW}=\alpha(\xi)m^{tSW}(\xi_{e})$, where $\xi_{e}=\xi+\chi_{L}\cdot\alpha(\xi)\cdot m^{tSW}(\xi)$, and $\alpha$ is a linear function of $\xi$. We see a rather good agreement between the analytical expression (orange and red dashed line, for $\chi_{L}=2.4$ and $\chi_{L}=4.8$, respectively) and the simulated data (orange and red square symbols, for $\chi_{L}=2.4$ and $\chi_{L}=4.8$, respectively). The results shown in Fig. 5 demonstrate the utility of the tSW approach in the study of interacting systems, and hint at the promise of similar highly scalable methods to open new horizons in the study of magnetic soft matter.
@@
In summary, the tSW model is a reasonably accurate approach to introduce internal magnetization dynamics for $\sigma\geq 5$ nanoparticles via Néel relaxation, that can reproduce the correct equilibrium in this $\sigma$ range. Having established that, the discussion can proceed to the study of dynamics, which is the type of study where the tSW model can benefit researchers interested in long timescales and bulk-sized systems.
@@ II.2 Dynamic Properties
One of the key collective properties of any magnetic soft matter system is its characteristic magnetic relaxation time. This is especially relevant for systems designed for heating applications. The magnetic relaxation time can be assessed by calculating the dynamic susceptibility spectra and locating the maximum/maxima of the susceptibility’s imaginary part.
Figure 7: Dynamic susceptibilities as a function of $\omega\tau_{D}$, where $\tau_{D}/\tau_{B}=0.02$, for a non-interacting (i) solid superparamagnet ((a) real part; (c) imaginary part), where the analytical tSW results (dashed lines) and simulations (symbols) are compared to the continuum-level description (solid lines); ferrofluid ((b) real part; (d) imaginary part), where the analytical tSW results (dashed lines) and simulations (symbols) are compared to the continuum-level description (solid lines). The analytical tSW corresponds to Eq. 33, where the input relaxation time is either the Néel time, or the effective time $\left(\tau^{-1}_{N}+\tau_{B}^{-1}\right)^{-1}$ (solid superparamagnet and ferrofluid, respectively) . The continuum-level description corresponds to Eq. 27 where the relaxation times $\tau_{\parallel}$ and $\tau_{\perp}$ are given either by Eq. 31 or Eq. 32 (solid superparamagnet and ferrofluid, respectively). The color bar indicates the anisotropy parameter $\sigma$ values explored.
Let us consider an ensemble of immobilized, non-interacting, randomly oriented particles with uniaxial magnetic anisotropy. In a harmonically oscillating field $H=H_{0}\cos\omega t$, where $t$ is the time and $\omega$ is the oscillation frequency, the magnetization can generally be represented as a series:
$m(t)=\sum_{k=1}^{\infty}\left(m^{\prime}_{k}(\omega)\cos k\omega t+m^{\prime\prime}_{k}(\omega)\sin k\omega t\right),$ (23)
@@
$\chi^{\prime}(\omega)=m^{\prime}_{1}(\omega)/H_{0},\>\>\chi^{\prime\prime}(\omega)=m^{\prime\prime}_{1}(\omega)/H_{0}.$ (25)
@@
$\chi=\left(\frac{1}{3}\chi_{\parallel}+\frac{2}{3}\chi_{\perp}\right),$ (27)
where, within the continuous approach and without gyromagnetic effects [61]:
$\chi_{\parallel}(\omega)$ $\simeq\frac{\chi_{\parallel}(0)}{1-i\omega\tau_{\parallel}},\>\>\>\chi_{\parallel}(0)=\chi_{L}(1+2S),$ (28)
$\chi_{\perp}(\omega)$ $\simeq\frac{\chi_{\perp}(0)}{1-i\omega\tau_{\perp}},\>\>\>\chi_{\perp}(0)=\chi_{L}(1-S),$
the order parameter $S$ is given by:
@@
$D(x)=e^{-x^{2}}\int_{0}^{x}e^{t^{2}}\,\text{d}t.$ (30)
This relation must be true for both the static and dynamic case. For immobilized particles:
$\tau_{\parallel}$ $=\tau_{D}\frac{e^{\sigma}-1}{2\sigma}\left[\frac{1}{1+1/\sigma}\sqrt{\frac{\sigma}{\pi}}+2^{-\sigma-1}\right]^{-1},$ (31)
$\tau_{\perp}$ $=2\tau_{D}\frac{1-S}{2+S},$
In the case where particles are not immobilized, it is sufficient to replace relaxation times with effective values [50]:
$\tau^{*}_{\parallel}=\left(\tau^{-1}_{\parallel}+\tau_{B}^{-1}\right)^{-1},\\ \tau^{*}_{\perp}=\left(\tau^{-1}_{\perp}+\tau_{B}^{-1}\right)^{-1}.$ (32)
It is important to note that field-dependent relaxation should also be scrutinized to fully appreciate the implications from the coupling of the magnetic relaxation parallel and perpendicular to the easy axis. As is, this remains an open question for study. In the tSW model, the relaxation time $\tau_{\parallel}$ is an input parameter, which means that the dynamics of tSW parallel to the easy axis must be the same as in the continuous case. The orthogonal relaxation time $\tau_{\perp}$ is instantaneous. Therefore, the dynamic susceptibility of an ensemble of frozen, uniaxial magnetic nanoparticles with a random distribution of anisotropy axis is given by:
@@
Once again, for a liquid, it is sufficient to introduce effective relaxation times: $\tau_{\parallel}\rightarrow\tau^{*}_{\parallel}$. Note that the orthogonal magnetic response (which is instantaneous and does not have any fluctuations associated with it) will not make any contribution to the zero-field magnetization autocorrelation function. Therefore, the calculated susceptibilities from the autocorrelation function when using the tSW method will be incorrect. The simulated susceptibilities have been calculated by applying a weak oscillating magnetic field (maximum amplitude corresponding to $\xi=0.1$) for a span of frequencies, in order to be able to determine the complex susceptibilities from the time evolution of the total magnetization vector (see Eq.24). Similarly, the static susceptibilities should be estimated from the initial slope in the magnetization profiles.
@@
Looking at Fig. 7, we can see that the tSW model reproduces the correct dynamics in both the solid superparamagnet and non-interacting ferrofluid systems, as quantified by the initial dynamic susceptibilities. Of course, in the initial susceptibility regime, the tSW model incorporated dipole moment fluctuations only along the anisotropy axis associated with $\tau_{\parallel}$, which in this case is equivalent to $\tau_{N}$. Since this is an input parameter for the model, the tSW can confidently reproduce the correct internal dynamics for the lowest frequency relaxation process ($\tau_{\parallel}$). It is apparent that Néel relaxation is very relevant even for $\sigma=10$ particles, where the dynamic susceptibility is clearly distinct from the Debye limit shown with dots. Recall that the maximum of the Debye limit is at $\omega\tau_{d}=0.01$, which is set by the model as $\tau_{D}/\tau_{N}=0.01$. The high frequency peaks, resolved in the continuum description, corresponding to $\tau_{\perp}$, cannot be captured in the tSW, simply due to the fact that the only intra-well states available to the dipole moment in the tSW model are the minima of the magnetic energy. In other words, there are no associated fluctuations perpendicular to the anisotropy axis in the $H\rightarrow 0$ regime. Outside of the initial susceptibility regime, this is not an issue as the dipole moment instantaneously follows the local energy minima, as obtained in the classical SW model, where there is always a relative angle the dipole moment takes with respect to the anisotropy axis. However, within the range of applicability we suggest ($\sigma>5$), these higher frequency processes are distinct enough that they could be averaged out while still resolving the correct dynamics. Having said that, it should also be clear that the in-field susceptibility contribution perpendicular to the anisotropy axis is frequency independent.
@@
As suggested at the beginning of this section, the position(s) of the maxima of the imaginary part of the initial susceptibility indicate the optimal field frequencies for heating applications. Fig. 7 clearly shows that if one needs to optimize the frequency of an applied magnetic field for hyperthermia depending on the particle size and material, using a point dipole approximation can be very misleading, as even for rather high anisotropy energies, the shape and the imaginary part of the susceptibility spectra drastically differ from that of a purely Brownian, or purely Néel relaxing systems.
@@
In summary, we can see that the tSW approach can be used effectively to incorporate Néel relaxation and simulate the dynamics associated with that relaxation process for magnetic nanoparticles with $\sigma\geq 5$. The relative importance of the higher-frequency mechanisms not considered in the tSW approach has to be determined on a case-by-case basis. The tSW approach could enable new simulation studies in cases where such mechanisms are not crucial. The price to pay in terms of accuracy is slight, compared to the simulation scale compromises that would be unavoidable with an approach that reproduces also the higher frequency relaxation processes. A major use case of the tSW model would be to simulate systems where $\tau_{B}$ is high. In this case, a Brownian dynamics approach to thermalize each particle might be impractical. A good example here would be multicore or composite magnetic colloids, where the nanoparticles are embedded in a solid matrix [79, 80, 77].